OpenStudy (anonymous):

write the equation of the line passing through the point (2,-5) parallel to y=7?

OpenStudy (anonymous):

...did you graph it?

OpenStudy (nincompoop):

With or without graph, this should not be difficult to solve.

OpenStudy (anonymous):

Y = mx +c you know m find c by putting values (2,-5) in the equation

OpenStudy (anonymous):

true im just saying the graph would give you a visual

OpenStudy (nincompoop):

When it says parallel to a particular line or slope, it just means that slope of the equation you are solving is the same as the given

OpenStudy (anonymous):

Let @VicVic decided :)

OpenStudy (anonymous):

Without a graph.

OpenStudy (anonymous):

K can you solve it from heere vic?

OpenStudy (anonymous):

No I am still lost.

OpenStudy (nincompoop):

And every time you are looking for the equation of a line, it is always good to prepare the y-intercept form Y= mx+b Where m is your slope and b is the y-intercept, which has the coordinate (0, b)

OpenStudy (anonymous):

I think Let me draw the graph as well 

OpenStudy (anonymous):

I know the point-slope is y-y1= m(x-x1). @nincompoop

OpenStudy (nincompoop):

Yeh

OpenStudy (anonymous):

So again which formula do i use to solve this problem?

OpenStudy (nincompoop):

So that means your line of reference has a zero slope y=m(0)+7

OpenStudy (anonymous):

Noooooo I don't understand what is the answer?

OpenStudy (nincompoop):

This also means that the one you are solving has a zero slope.

OpenStudy (nincompoop):

We don't give out answers, coocoobird

OpenStudy (anonymous):

What is coocoobird?

OpenStudy (anonymous):

take the slope same as y=7 and choose c in (y-y1)=m(x-x1) just insert the points

OpenStudy (nincompoop):

Using the point-slope form with zero slope it would look like this Y-y1 = 0 (x-x1) Solve the rest, scoochielooch

OpenStudy (nincompoop):

IF you don't know how solve from this point on, then you've got a bigger problem

OpenStudy (nincompoop):

I'll tie you a technique later when you grasp the whole concept so your life will be easier

OpenStudy (nincompoop):

Do you understand, coocoovic?

OpenStudy (anonymous):

Do not call me that. I know exactly what it means. Thank you but your help is no longer needed. @nincompoop